A retailer has some skirts that cost $25 each. She wants to sell them at a profit of 50% of the selling price. What price should she charge for the skirts?
$


Found 2 solutions by jorel555, stanbon:
Answer by jorel555(1290)

$25 + 50%=25+12.50=$37.50 selling price.

Answer by stanbon(75887)

A retailer has some skirts that cost $25 each. She wants to sell them at a profit of 50% of the selling price. What price should she charge for the skirts?
---------------------
Ans: 25 + 0.50*25 = 1.5*25 = $37.50
----------------------
Cheers,
Stan H.
-------------